# Optimal Computation in Anonymous Dynamic Networks

### v7

@@ 7 Conclusions
We introduced the novel concept of history tree and used it as our main investigation technique to study computation in anonymous dynamic networks. History trees are a powerful tool that completely and naturally captures the concept of symmetry and indistinguishability among processes. We have demonstrated the effectiveness of our methods by optimally solving a wide class of fundamental problems, and we believe that our techniques will find numerous applications in other settings, as well.
@@
We have shown that anonymous processes in $\tau$-union-connected dynamic networks can compute all the multiset-based functions and no other functions, provided that the network contains a known number of leaders $\ell\geq 1$. If there are no leaders or the number of leaders is unknown, the class of computable functions reduces to the frequency-based functions. We have also identified the Input Frequency function and the Input Multiset function as the complete problems for each class. Notably, the network’s dynamic disconnectivity $\tau$ does not affect the computability of functions, but only makes computation proportionally slower.
@@
Moreover, we gave efficient stabilizing and terminating algorithms for computing all the aforementioned functions. Some of our algorithms make assumptions on the processes’ a-priori knowledge about the network; we proved that such assumptions are actually necessary. All our algorithms have optimal linear running times in terms of $\tau$ and the size of the network $n$.
@@
In one case, there is still a small gap in terms of the number of leaders $\ell$. Namely, for terminating computations with $\ell\geq 1$ leaders, we have a lower bound of roughly $\tau(2n-\ell)$ rounds (Proposition 6.8) and an upper bound of roughly $\tau(\ell^{2}+\ell+1)n$ rounds (Theorem 5.2). Although these bounds asymptotically match if the number of leaders $\ell$ is constant (which is a realistic assumption in most applications), optimizing them with respect to $\ell$ is left as an open problem.
@@
It is worth noting that for stabilizing computation (i.e., when explicit termination is not required) in networks with a constant number of leaders, our lower and upper bounds are essentially $2\tau n$ rounds; hence, in these case we were able to optimize the multiplicative constant, as well. As for terminating computation with a unique leader, we have a lower bound of $2\tau n$ rounds and an upper bound of $3\tau n$ rounds. Although we are still unable to completely close this gap, we emphasize that our findings demonstrate the practical feasibility of general computations in anonymous dynamic networks with a unique leader, which was a major open problem in this research area prior to our work.
@@
Observe that our stabilizing algorithms use an unbounded amount of memory, as processes keep adding nodes to their views at every round. This can be avoided if the dynamic disconnectivity $\tau$ (as well as an upper bound on $n$, in the case of a leaderless network) is known: In this case, processes can run the stabilizing and the terminating version of the relevant algorithm in parallel, and stop adding nodes to their views when the terminating algorithm halts.
@@
Our algorithms require processes to send each other explicit representations of their history trees, which have roughly cubic size in the worst case. It would be interesting to develop algorithms that only send messages of logarithmic size, possibly with a trade-off in terms of running time. We are currently able to do so for leaderless networks and networks with a unique leader, but not for networks with more than one leader [24].
@@
Finally, we wonder if our results hold more generally for networks where communications are not necessarily synchronous. We conjecture that our algorithms can be generalized to networks where messages may be delayed by a bounded number of rounds or processes may be inactive for some rounds (provided that a “global fairness” condition is met).

### v8

@@ 7 Conclusions
We introduced the novel concept of history tree and used it as our main investigation technique to study computation in anonymous dynamic networks, modeled as sequences of multigraphs. History trees are a powerful tool that completely and naturally captures the concept of symmetry and indistinguishability among agents. In fact, the history tree of a (static or dynamic) network encodes all the information that can be extracted by the agents in the network (Theorem 3.1). We have demonstrated the effectiveness of our methods by optimally solving a wide class of fundamental problems (Table 1), and we believe that our techniques will find numerous applications in other settings, as well.
@@
We have shown that anonymous agents in $\tau$-union-connected dynamic networks can compute all the multiset-based functions and no other functions, provided that the network contains a known number of leaders $\ell\geq 1$. If there are no leaders or the number of leaders is unknown, the class of computable functions reduces to the frequency-based functions. We have also identified the Input Frequency function and the Input Multiset function as the complete problems for each class. Notably, the network’s dynamic disconnectivity $\tau$ does not affect the computability of functions, but only makes computation proportionally slower.
@@
Moreover, we gave efficient stabilizing and terminating algorithms for computing all the aforementioned functions. Some of our algorithms make assumptions on the agents’ a-priori knowledge about the network; we proved that none of these assumptions can be removed. All our algorithms have optimal linear running times in terms of $\tau$ and the size of the network $n$.
@@
In one case, there is still a small gap in terms of the number of leaders $\ell$. Namely, for terminating computations with $\ell\geq 1$ leaders, we have a lower bound of roughly $\tau(2n-\ell)$ rounds (Proposition 6.8) and an upper bound of roughly $\tau(\ell^{2}+\ell+1)n$ rounds (Theorem 5.2). Although these bounds asymptotically match if the number of leaders $\ell$ is a constant (which is a realistic assumption in most applications), optimizing them with respect to $\ell$ is left as an open problem.
@@
It is worth noting that for stabilizing computation (i.e., when explicit termination is not required) in networks with a fixed number of leaders, our lower and upper bounds are essentially $2\tau n$ rounds; hence, in this case we were able to optimize the multiplicative constant, as well. As for terminating computation with a unique leader, we have a lower bound of $2\tau n$ rounds and an upper bound of $3\tau n$ rounds. Although we are still unable to completely close this gap, we emphasize that our findings demonstrate the practical feasibility of general computations in anonymous dynamic networks with a unique leader, which was a major open problem in this research area prior to our work.
@@
Observe that our stabilizing algorithms use an unbounded amount of memory, as agents keep adding nodes to their vistas at every round. This can be avoided if the dynamic disconnectivity $\tau$ (as well as an upper bound on $n$, in the case of a leaderless network) is known: in this case, agents can run the stabilizing and the terminating version of the relevant algorithm in parallel, and stop adding nodes to their vistas when the terminating algorithm halts.
@@
Our algorithms require agents to send each other explicit representations of their history trees, which have roughly cubic size in the worst case. It would be interesting to develop algorithms that only send messages of logarithmic size, possibly with a trade-off in terms of running time. We are currently able to do so for leaderless networks and networks with a unique leader, but not for networks with more than one leader [28].
@@
Finally, we wonder if our results hold more generally for networks where communications are not necessarily synchronous. We conjecture that our algorithms can be generalized to networks where messages may be delayed by a bounded number of rounds or agents may be inactive for some rounds (provided that a “global fairness” condition is met).
Acknowledgments. The authors would like to thank the anonymous reviewers of the Journal of the ACM for corrections and suggestions that greatly improved the readability of this paper. G. A. Di Luna was supported in part by the Ministero dell’Università e della Ricerca (MUR) National Recovery and Resilience Plan funded by the European Union-NextGenerationEU through the Project SEcurity and RIghts In the CyberSpace (SERICS) under Grant PE00000014 and project ATENEO Sapienza RM1241911201548A. G. Viglietta was partially funded by the JSPS KAKENHI Grant number 23K10985 and the University of Aizu Competitive Research Fund.